OpenStudy (anonymous):

(cos Θ − cos Θ)2 + (cos Θ + cos Θ)2 sin2 Θ 4cos2 Θ 8 cos2 Θ

OpenStudy (anonymous):

@JoannaBlackwelder Do you know how to solve?

OpenStudy (anonymous):

@MathStudent420

OpenStudy (xapproachesinfinity):

the first one cancels costheta -costheta=0 \[(\cos \theta +\cos \theta)^2=(2\cos \theta)^2=4\cos^2 \theta\]

OpenStudy (anonymous):

That is what I had! Ugh feels great to be right. Canyou help with one more?

OpenStudy (anonymous):

f(x) = 2 sin(2x − π) + 2 g(x) graph of a quadratic with points at 1, 2 and 3, negative 2 and 5, 2 h(x) x y −2 10 −1 7 0 5 1 3 2 5 3 7 4 10 Which function has the smallest minimum? All three functions have the same minimum f(x) g(x) h(x)

OpenStudy (xapproachesinfinity):

hmm the first has minimum whenever 2sin(2x-pi)+2=-1 the second do you mean points (1,2) (3,-2) (5,2)? the third has minimum at x=1 which is 3

OpenStudy (anonymous):

It's a parabola on a graph.

OpenStudy (xapproachesinfinity):

first you would have to solve to find where exactly the min occurs

OpenStudy (anonymous):

The bottom of the parabola which is at -2

OpenStudy (xapproachesinfinity):

-2 would be the minimum the way you wrote it is not good that's why i'm asking you should write ordered pairs (x,y) like this

OpenStudy (anonymous):

It is a chart, hold on let me send u a picture of it

OpenStudy (xapproachesinfinity):

solve the first one to get the min

OpenStudy (xapproachesinfinity):

\[2\sin (2x-\pi)+2=-1 \]

OpenStudy (anonymous):

That is the graph they give

OpenStudy (xapproachesinfinity):

yeah the min for that is -2

OpenStudy (anonymous):

and for h(x) it is -2 as well

OpenStudy (xapproachesinfinity):

no it is 3 according to that table

OpenStudy (anonymous):

oh ok the smallest is g(x)???? Because we have -1,-2, and 3

OpenStudy (xapproachesinfinity):

hmm no the first is not -1 you have to solve to see where that occurs

OpenStudy (xapproachesinfinity):

do you know how to solve such equations?

OpenStudy (anonymous):

no I dont think so

OpenStudy (xapproachesinfinity):

is this calculus based question + how come you don't know how to solve equations that involve trigonometry? it is part of algebra 1 or so

OpenStudy (anonymous):

Algebra 2

OpenStudy (anonymous):

I forgot how to solve it!

OpenStudy (xapproachesinfinity):

will look for how to do it then

OpenStudy (xapproachesinfinity):

well*1

OpenStudy (xapproachesinfinity):

i believe you are given in notes how to deal with sins and cosine and how to get min and max

OpenStudy (anonymous):

We are. Is the answer on the graph -3.3?

OpenStudy (xapproachesinfinity):

given \[f(x)=A\sin(Bx-C)+D\] the min is D-|A| in our example \[f(x)=2\sin(2x-\pi)+2\] the min is 2-2=0

OpenStudy (anonymous):

So g(x) is still the answer?

OpenStudy (xapproachesinfinity):

yeah

OpenStudy (anonymous):

Thanks so much for your help

OpenStudy (xapproachesinfinity):

np